OpenStudy (anonymous):

Let f(x) = 3x^2 + 3x − 8. Find and simplify the following. f(a) + f(2)

OpenStudy (anonymous):

Okay, so for this one you'd plug in a everywhere you see x which would give you 3a^2+3a-8 and then 2 for everywhere you see x which would give you 3(2)^2+3(2)-8. So you'd end up having 3a^2+3a-8+12+6-8

OpenStudy (anonymous):

i have no clue about that. have u also tried googling it?

OpenStudy (anonymous):

And you just simplify from there! Giving you: 3a^2+3a+2 (because (-8+-8)=-16 and (12+6)=18 so, 18-16=2) Then you just "bust the b" which i can draw for you

OpenStudy (anonymous):

ah okay thank you!!

OpenStudy (anonymous):

Actually, I just tried it, and I think that is the simplified form!

OpenStudy (anonymous):

yep your right